OpenStudy (anonymous):

PLEASE HELP!!!!!!!!!!!!!!!!!!!!!!!!!!!!! Find the solution(s) to the system. Round answers to the nearest hundredth, if necessary. y=-0.5x^2+6 y=3.25

OpenStudy (anonymous):

y=\[-0.5x^2+6=3.25\]

OpenStudy (anonymous):

x^2=5.5

OpenStudy (theeric):

But, would you round up, @Krishnadas ?

OpenStudy (anonymous):

isnt this rounded to the 100th?

OpenStudy (theeric):

I got \(\sqrt{5.5}\approx 2.3452078799117147772828150567722\) And since the thousandths place is 5, you add one-one hundredth.

OpenStudy (anonymous):

thats how its done?In physics 2.345=2.34

OpenStudy (anonymous):

idk..

OpenStudy (theeric):

And, I don't know when this is taught, but it would be \(\pm\).

OpenStudy (theeric):

For real numbers.

OpenStudy (theeric):

Huh, I wasn't taught that...

OpenStudy (anonymous):

hmm.ok.ok

OpenStudy (anonymous):

can you round it up?

OpenStudy (theeric):

I was always taught if the number was 5 or higher you round up. If the number is like \(.2345000\), forever zeroes, then the number is exactly between \(2.34\) and \(2.35\). If it is any more, even one-millionth, then it is closer to \(2.35\).

OpenStudy (theeric):

Even one-quintillionth. Which I guess is \(10^{-18}\), which I never knew but makes sense.

OpenStudy (anonymous):

so..what should be the final answer

OpenStudy (theeric):

I guess it depends on which method you use. I've learned my way since I was young, and it makes more sense to me.

OpenStudy (anonymous):

2.35

OpenStudy (anonymous):

@theEric there is this rounding off calculator in the internet which gave 2.35

OpenStudy (theeric):

http://www.mathsisfun.com/numbers/rounding-methods.html See the "common way" I like the symmetric way! :) So 2.35.